if in a triangle tanA/2,tanB/2,tanC/2 are in H.P then the value of cotB/2

Grade:11

1 Answers

Harsha
6 years ago
Cot(A/2),Cot(B/2),Cot(C/2) are in AP2Cot(B/2)=Cot(A/2)+Cot(C/2)3Cot(B/2)=Cot(A/2)+Cot (B/2)+Cot (C/2)3Cot(B/2)={a*2+b*2+c*2}/4∆ThereforeCot(B/2)={a*2+b*2+c*2}/12∆